fhjfjfgjgfjyhtjgfyj:

Which statement best describes (5 × 7) × 3 = 5 × (7 × 3)?

RaeRae:

Hey! What are your options?

RaeRae:

and What do you think it is-

snowflake0531:

The associative property states \( a \times (b \times c) = (a \times b) \times c\)